# SpeechCap: Leveraging Playful Impact Captions to Facilitate Interpersonal Communication in Social Virtual Reality

### v1

@@ 4.3. Design Goals
Based on the results of the co-design, we propose the following design goals to implement a proof-of-concept tool to support interpersonal communication in social VR through impact captions.
4.3.1. G1: Generating impact caption in time
@@ 5. SpeechCap: System Design and Implementation
5.1. Technical Pipeline: Algorithms, Software, and Hardware Implementations
The core technical pipeline of SpeechCap (Figure 4) consists of three main software modules: (1) a Voice Interface that transcribes speech to text in real time and stores the transcribed text for subsequent processing; (2) a Text Processor that decides which words should be made as impact captions and the appearance of each impact caption to be rendered in VR based on the analysis of user speech. (3) a VR Application that provides a multi-user VR environment for users to realistically enjoy communication using interactive impact captions.
Figure 4. Technical Pipeline of SpeechCap. The system consists of Voice Interface, Text Processor, and VR Application. The Voice Interface processes users’ real-time voice input and stores the transcribed texts. The VR Application keeps polling the Text Processor for fetching upcoming impact captions, driving the Text Processor to convert transcribed texts into impact captions with customized visual design.
@@ 5.1.1. Voice Interface
We implemented a Python module with basic real-time speech-to-text capability using OpenAI’s Whisper  [^5] https://github.com/openai/whisper  (Radford et al., 2023) as the core model for speech recognition. As a state-of-the-art AI model designed for speech recognition tasks, Whisper’s accuracy is close to that of professional human transcribers  (Radford et al., 2023), and its speed is above 10 words per second when using the minimized pruned model (i.e., “tiny.en”) for transcribing English on an ordinary hardware platform  (Haz et al., 2023).
@@
Since Whisper was not originally designed for real-time transcribing, we employed a self-developed multi-thread algorithm to approximately process voice input in a real-time manner. When starting, the algorithm recurrently chunks the input voice stream into fragments with a fixed duration (e.g. 1 second by default). Once a fragment is created, the algorithm would allocate a thread to call Whisper for transcribing, and then save the text result with a sequence number. When running the algorithm, the sequence number starts from 0 and increases by 1 each time for a newly processed fragment. The sequence number mechanism ensures that the downstream Text Processor and VR Application can always keep alignment with the original speech when they process texts and render impact captions in VR.
@@
In terms of accuracy, our method could be slightly lower than the core model because we rely on the multi-threading strategy to transcribe the chunks of speech input one-by-one. Thus, a single word could possibly be separated into two successive chunks and then fail to be recognized, even if the chance is small. For speeding up, Whisper provides a sort of pruned model with smaller sizes to reduce computational consumptions. And in our algorithm, the maximum number of simultaneous working threads is configurable. More available threads can accelerate transcribing while requiring more computational power. Ideally, the delay between a word being recorded and transcribed can be calculated by summing up the chunk duration (e.g. 1 second), Whisper processing time, and the time of waiting for an available thread. Particularly, in our user study, we adopted the “tiny.en” model of Whisper (with the minimum number of parameters and specified for English) and a maximum of 4 threads to maintain an acceptable transcribing speed for running the Voice Interface on Apple MacBook Pro with the M1 Pro chipset.
@@
Our real-time speech recognition algorithm is scalable in terms of both accuracy and efficiency. By employing Whisper models with different sizes or other SOTA speech recognition models, both dimensions could possibly be improved, although the trade-off between accuracy and efficiency always holds. Except for the core model, by configuring different chunk durations and the number of available threads (i.e. thread pool size), the efficiency is optimizable. In addition, advanced hardware resources other than laptops can bring significant improvements to the performance of the Voice Interface module.
@@ 5.1.2. Text Processor
To decide which words should be made as impact captions and the specific visual design of each individual caption, we developed the Text Processor module to analyze transcribed speech using NLTK  [^6] https://www.nltk.org/ and Pydub  [^7] https://pydub.com/. Text Processor and upper-stream Voice Interface are connected by a shared database (SQLite) in which transcribed texts are stored by Voice Interface. And to respond to the needs of the downstream VR Application, we encapsulate the Text Processor in a back-end web server using the Flask framework  [^8] https://flask.palletsprojects.com/. The server is responsible for providing processed data (i.e., the content and visual design of impact captions) for the VR Application through HTTP APIs.
@@ 5.1.3. VR Application
Nevertheless, the replication mechanism does not force every “client” to strictly have the same appearance all the time, since each individual “client” always reserves full control of the objects in its own simulation. This allows us to make local modifications to the objects before they are finally rendered on each individual “client”. In this way, we adjust the rotation parameters of the impact captions on each “client” HMD, making the impact captions always face towards the HMD’s owner player in a readable angle (i.e., the replicants of the “same” impact caption are rendered to face different directions in different “clients”).
@@ 5.1.4. Hardware Settings
Hardware settings include a laptop, a wireless microphone, and a VR head-mounted device with controllers for a single user. Specifically, we employed an Apple MacBook Pro (with an M1 Pro chipset) laptop for running the Voice Interface and Text Processor modules with the associated database and web server, and a Bluetooth wireless microphone connected to the laptop for collecting the user’s voice input. Meta Quest 2 with controllers is used to run the SpeechCap VR Application.
@@ 5.2. Speech-driven Impact Caption Generation
In SpeechCap, we adopted a rule-based method to determine the specific visual appearance for impact captions on each dimension of the design space. Our method takes verbal content, affective factors (e.g., valence) and prosodic factors (e.g., loudness) extracted from the speech to compute the specific visual design of each impact caption (Figure 5), making impact captions appropriate for the corresponding conversations.
@@ 5.3. Communication with Interactive Impact Captions in VR
SpeechCap allows users to have conversations in a shared VR space with the support of interactive impact captions. In SpeechCap, impact captions can only be triggered by speech input. When a particular word is mentioned in the speech, SpeechCap will create a corresponding impact caption at a position roughly in front of the speaker’s avatar at the height of the chest so that users can easily grab the captions. To avoid overlapping, the position is slightly randomized each time for a newly incoming caption (subsubsection 5.3.4).
The visual appearance of each caption instance is determined by the semantics and affective factors of the corresponding speech fragment (subsection 5.2). During a conversation, both the speaker and the listener can see and play with the captions, making their communication process visible, interactive, and playful beyond the traditional voice-only experience.
Figure 6. Interactions with Impact Caption in SpeechCap. (A) Grab to hold and place a caption to a new position; (B) Stretch with two hands to resize an caption; (C) Attach an caption to (another player’s) head; (D) Shoot an caption to trigger the Explosion effect.
@@ 5.3.1. Fundamental Interactions
SpeechCap enables users to play with impact captions in VR intuitively using a set of embodied actions (G3), which achieves fundamental interactions between user and the captions. Grab (Figure 6 A.) is an action that allows users to hold a caption with their hands using controllers, change its orientation, and place it elsewhere just like taking a physical item in the real world. When grabbing an impact caption, users can further Shake the caption to make it start shivering and blinking. When both hands grab the same impact caption simultaneously, the caption can be Stretched (Figure 6 B.) for resizing. By resizing, users may create huge captions larger than their avatars, or tiny captions like a bullet. The scalability of size extends the usage of impact captions beyond speech-driven subtitles. Users may combine captions of different sizes with other actions to achieve interactions with others. For example, making a huge caption and shaking it to trigger the blinking effects to enhance the visual impact to emphasize the key points of conversation.
@@ 5.3.2. Decorate Avatars for Self Expression
Avatar plays an important role in social VR that supports self-presentation and reflects users’ emotions and personalities  (Sykownik et al., 2022; Freeman and Acena, 2021). In SpeechCap, users can Attach (Figure 6 C.) an impact caption to the top of their virtual head by tapping or lightly touching the head while holding the caption. Once attached, the caption could be fixed continuously like a “tag” or “hat”. This feature is inspired by the typical usage of impact captions in TV shows where impact captions are used to tag the characters, revealing their emotional status. Besides the head, SpeechCap also allows users to attach an impact caption to the body part of avatars, in which the attached caption will start to periodically revolve around the avatar, achieving the surrounding motion effect in the design space.
@@ 5.3.3. Mediating Interpersonal Interactions
In SpeechCap, impact captions can react to the actions and avatars of multiple users so that they can be used as a medium to facilitate the interactions between the users. When thrown or shot by users, impact captions start flying in the virtual space. Once the flying impact caption crashes with another impact caption or avatar, an explosion effect will be triggered (Figure 6 D.), in which several duplications of the flying caption will be generated at the crashing point and emit outwards fast, like the explosion of fireworks. Such a dynamic effect may be used as a non-verbal cue to attract attention and add more fun to the social VR experience.
@@ 5.3.4. Reducing Visual Clutter and Overlapping
To maintain the readability and clarity of floating impact captions in conversations (G2), SpeechCap introduces a “time-to-live (TTL)” mechanism to softly control the number of captions. The “TTL” mechanism allows each caption to live for 5 seconds once generated. If users do not interact with a newly generated impact caption within the time period, the caption will automatically disappear afterwards. Otherwise, if the user wants to keep the impact caption for any purpose, the simplest action to do so is to touch the caption or to grab it at least once before it disappears. Once the caption has interactions with a user, it will no longer be constrained by the living time limitation until users intentionally delete it by holding it and clicking a button on the VR controller (i.e., “X” button for the left-hand controller and “A” button for the right-hand controller).
Additionally, to avoid multiple impact captions overlapping with each other when a user keeps speaking, SpeechCap randomizes the positions for spawning. Specifically, the exact spawn position of a caption is calculated by a fixed position plus random offsets on $x$, $y$, and $z$ dimensions. As a result, this strategy makes multiple impact captions look like a “word cloud” emerging in front of the speaker’s avatar.
@@ 6.1. Making Social VR Accessible to Deaf and Hard-of-Hearing (DHH) People
Recent research revealed that current VR applications fail to provide sufficient accessibility support for deaf and hard-of-hearing (DHH) people to neither experience immersive digital content nor conduct remote communication and socialization  (Jain et al., 2021; Borna et al., 2024). Captioning systems, known as useful tools for making content accessible in other digital media such as videos, are still under-explored in VR contexts  (Kim et al., 2023; de Lacerda Pataca et al., 2023, 2024). SpeechCap not only have similar functionalities as previous systems that turn speech into visible captions for DHH people  (Kim et al., 2023; de Lacerda Pataca et al., 2023; Li et al., 2022), but also extend the idea by incorporating multiple visual cues and adding interactivity to convey rich non-verbal information (i.e., valence, excitement, etc.) in virtual reality.
For example, when participating in a conversation with someone who is talking about a vacation travel plan in social VR with SpeechCap, a DHH user can perceive the speaker’s positive mood through the captions in a bright color with a laughing emoji. An additional shivering motion can notify the user that the speaker is also really excited about the trip. By dragging and placing impact captions of the names of the must-visit spots mentioned in speech, the speaker can demonstrate the plan with clarity. Further, when the speaker picks up a caption containing “volcano” and talks about her passion for hiking and exploration, the impact caption with a corresponding volcano icon can give clear reference to the ongoing speech. Overall, informative captions, engaging visuals, and interactive playfulness provided by SpeechCap can enlarge the space for DHH individuals to communicate with other people and the virtual world in social VR.
@@ 6.2. Enhancing Interactions for Teaching and Learning in Social VR
VR has shown a great potential in education for the increased social presence it provides in remote learning, and the access to a wide range of learning contexts that may not be accessible in reality  (Thanyadit et al., 2022; Peng et al., 2022; Jensen and Konradsen, 2018). However, it is still far from being a commonly used tool due to the challenges of creating educational content that is easy for reviewing, providing inclusive functionalities, and supporting collaborative learning  (Jin et al., 2022). In a virtual lecture, SpeechCap allows instructors to create instructional cues (e.g., notes) by placing impact captions of keywords in the space. The shared interactive captions can further act as a medium for the interactions between the instructor and the students, facilitating learning activities like question answering or hands-on demonstration.
@@ 6.3. Facilitating Engaging Live Streaming Experiences in Social VR
Recent research suggests that VR streamers face challenges in building emotional connections since their facial expressions can be obscured by VR headsets, making it difficult for viewers to perceive the emotions of the streamer directly  (Wu et al., 2023), while emotional connection is a key factor in live streaming  (Lu et al., 2018). SpeechCap can provide visualized emotions for streamers by detecting the expressed moods within the speech and adding appropriate emoticons, motions to relevant captions. SpeechCap can also enhance the communication between the streamer and the viewers by rendering the viewers’ reactions as interactive captions in the streamer’s virtual space, mediating the interactions between the streamer and the viewers. This could be a way to satisfy the demands of visible and spatialized objects in engaging live streaming experiences  (Wu et al., 2023; Lu et al., 2018).
@@ 7.1. Participants
We recruited 14 participants (6 female and 8 male, aged between 18 and 35) from local universities by advertising on social media. There is no overlap between this group of participants and the four experts who contributed to the previous co-design activity. In terms of social VR experiences, five participants (annotated P1 to P5) were novice users who had never used VR HMDs before, while the other nine participants (P6 to P14) self-reported that they have rich experiences in VR games, social VR features, and VR application developments.

### v3

@@ 4.3. Design Goals
Based on the qualitative findings from the co-design activity, we propose the following design goals to implement a proof-of-concept tool to support interpersonal communication in social VR through impact captions.
4.3.1. G1: Generating Impact Captions in Time
@@ 5. SpeechCap: System Design and Implementation
5.1. Technical Pipeline: Algorithms, Software, and Hardware
The technical pipeline of SpeechCap (Figure 4) consists of three main software modules namely Voice Interface, Text Processor, and VR Application. Voice Interface transcribes speech to text in real time and stores the transcribed text for subsequent processing. Text Processor determines which words should be made as impact captions and decide the appearance of captions based on the analysis of speech texts. VR Application provides VR environment for users to engage in communication through interactive impact captions.
Figure 4. SpeechCap System Overview. The system consists of three key modules: (A) Voice Interface that processes real-time voice input and stores the transcribed texts to a shared database, (B) Text Processor that distills transcribed texts into impact captions and decides the design of each caption, and (C) VR Application that keeps polling the Text Processor for fetching upcoming impact captions and renders the VR space. As for hardware settings, Voice Interface and Text Processor run on a laptop that is paired with a Bluetooth Microphone for collecting voice input. A local area network (LAN) is configured to support the connection between the laptop and VR headsets and among multiple sets of hardware devices for multiple users.
@@ 5.1.1. Voice Interface
We implemented a Python module with basic real-time speech-to-text capability using OpenAI’s Whisper  [^5] https://github.com/openai/whisper  (Radford et al., 2023) as the core model for speech recognition. As a state-of-the-art AI model designed for speech recognition tasks, Whisper’s accuracy is close to that of professional human transcribers  (Radford et al., 2023), and a speed exceeding 10 words per second when using the minimal pruned model (i.e., “tiny.en”) for transcribing English on an ordinary hardware platform  (Haz et al., 2023).
@@
Since Whisper was not originally designed for real-time transcribing, we employed a custom multi-thread algorithm to process voice input in near real time. At the start, the algorithm divides the input voice stream into fragments of fixed duration (e.g., 1 second by default). Once a fragment is created, the algorithm would allocate a thread to call Whisper for transcribing, and then save the text result with a sequence number. When running the algorithm, the sequence number starts from zero and increases by one each time for a newly processed fragment. The sequence number mechanism ensures that the downstream Text Processor and VR Application can always maintain alignment with the original speech when they process texts and render impact captions in VR.
@@
In terms of accuracy, our method may slightly underperform compared to the core model due to the use of a multi-threading strategy that processes speech chunks sequentially Thus, a single word might be split into two successive chunks and fail to be recognized, although this is unlikely. For speeding up, Whisper provides a sort of pruned model with smaller sizes to reduce computational consumptions. And in our algorithm, the maximum number of simultaneous working threads is configurable. More available threads can accelerate transcribing while requiring more computational power. Ideally, the delay between a word being recorded and transcribed can be calculated by summing up the chunk duration (e.g. 1 second), Whisper processing time, and the time of waiting for an available thread. Particularly, in our user study, we adopted the “tiny.en” model of Whisper (with the minimum number of parameters and specified for English) and a maximum of 4 threads to maintain an acceptable transcribing speed for running the Voice Interface on Apple MacBook Pro with the M1 Pro chipset.
@@
Our speech recognition algorithm is scalable in terms of both accuracy and efficiency. By employing Whisper models with different sizes or other SOTA speech recognition models, both dimensions could be improved, although the trade-off between accuracy and efficiency remains. Except for the core model, by configuring different chunk durations and the number of available threads (i.e. thread pool size), the efficiency is optimizable. In addition, advanced hardware resources other than laptops can bring significant improvements to the performance of the Voice Interface module.
@@ 5.1.2. Text Processor
To decide which words should be made as impact captions and the specific visual design of each individual caption, we developed the Text Processor module to analyze transcribed speech using NLTK  [^6] https://www.nltk.org/ and Pydub  [^7] https://pydub.com/. Text Processor and upper-stream Voice Interface are connected by a shared database (SQLite) in which transcribed texts are stored by Voice Interface. And to respond to the needs of the downstream VR Application, we encapsulate the Text Processor in a back-end web server using the Flask framework  [^8] https://flask.palletsprojects.com/. The server is responsible for providing processed data (i.e., the content and visual design of impact captions) for the VR Application through Web APIs.
@@ 5.1.3. VR Application
Nevertheless, the replication mechanism does not force every “client” to strictly have the same appearance all the time, since each individual “client” always reserves full control of the objects in its own simulation. This allows us to make local modifications to the objects before they are finally rendered on each individual “client”. In this way, we adjust the rotation parameters of the impact captions on each “client” HMD, making the impact captions always face towards the HMD’s owner player in a readable angle (i.e., the replicas of the “same” impact caption are rendered to face different directions in different “clients”).
@@ 5.1.4. Hardware Settings
Hardware settings include a laptop, a wireless Bluetooth microphone, and a VR head-mounted device with controllers for a single user (Figure 6). Specifically, we employed an Apple MacBook Pro (with an M1 Pro chipset) laptop for running the Voice Interface and Text Processor modules with the associated database and web server, and a Bluetooth wireless microphone connected to the laptop for collecting the user’s voice input. Meta Quest 2 with controllers is used to run the SpeechCap VR Application.
@@ 5.2. Speech-driven Impact Caption Generation
In SpeechCap, we used a rule-based method to determine the visual appearance of impact captions across each dimension of the design space. Our method takes verbal content, affective factors (e.g., valence) and prosodic factors (e.g., loudness) extracted from the speech to compute the visual design of each impact caption (Figure 5), ensuring that the captions align with the context of the conversation.
@@ 5.3. Communication with Interactive Impact Captions in VR
SpeechCap allows users to have conversations in a shared VR space with the support of interactive impact captions. In SpeechCap, impact captions are exclusively triggered by speech input. When a particular word (i.e., identified by the rules in Text Processor, subsubsection 5.1.2) is mentioned in the speech, SpeechCap will create a corresponding impact caption at the position roughly in front of the speaker’s avatar at the height of the chest so that users can easily grab the captions. To avoid overlapping, the position will be randomized each time for newly generated captions (subsubsection 5.3.4).
The visual appearance of each caption is determined by the semantics and affective factors of the corresponding speech (subsection 5.2). In a conversation, both the speaker and the listener can see and play with the captions, allowing the communication process to be visible, interactive, and playful beyond traditional voice-only experiences.
Figure 6. Interactions with Impact Caption. (A) Grabbing allows users to hold and place a caption to an arbitrary position; (B) Stretching needs two hands to resize an impact caption; (C) Attaching allows an impact caption to be attached on the head or body of the virtual avatar; (D) Shooting can eject an impact caption forward and trigger the Explosion effect when a collision occurs.
@@ 5.3.1. Fundamental Interactions
SpeechCap allows users to intuitively interact with impact captions in VR through a set of embodied actions (G3), enabling basic human-caption interactions. Grab (Figure 6 A) allows users to hold a caption using VR controllers, adjust its orientation, and move it to a new position, mimicking real-world interactions with a physical item. When grabbing an impact caption, users can Shake it to trigger shivering and blinking effects.
If both hands grab the same impact caption simultaneously, users can Stretch it to resize (Figure 6 B). By resizing, users may create huge captions larger than their avatars, or tiny captions like a bullet. The ability to adjust size broadens the functionality of impact captions, making them more than just speech-driven subtitles. Users can combine captions of different sizes with other actions to enhance interactions with others. For example, making a huge caption and shaking it to trigger the blinking effects to enhance the visual impact to emphasize the key points of conversation.
@@ 5.3.2. Decorate Avatars for Self Expression
Avatars play an important role in social VR that supports self-presentation and reflects users’ emotions and personalities  (Sykownik et al., 2022; Freeman and Acena, 2021). In SpeechCap, users can Attach an impact caption to their avatar’s head by tapping or touching it while holding the caption (Figure 6 C). Once attached, the caption remains fixed like a “tag” or “hat”. This feature is inspired by the typical usage of impact captions in TV shows where impact captions are used to tag the characters, revealing their emotional status. Besides the head, users can also attach an impact caption to a body part of the avatar, causing it to revolve around the avatar periodically and creating a circling motion effect.
@@ 5.3.3. Mediating Interpersonal Interactions
In SpeechCap, impact captions can react to user actions and avatars, serving as a medium to facilitate interactions. When thrown or shot by users, impact captions start flying in the virtual space. Once a flying impact caption collides with another caption or avatar, it triggers an explosion effect (Figure 6 D), generating multiple replicas that emit outward like fireworks. Such a dynamic effect may be used as a non-verbal cue to attract attention and add more fun to the social VR experience.
@@ 5.3.4. Reducing Visual Clutter and Overlapping
To maintain the readability and clarity of floating impact captions in conversations (G2), SpeechCap introduces a time-to-live (TTL) mechanism to manage the number of captions. The TTL mechanism allows each caption to live for 5 seconds once generated. If users do not interact with a newly generated impact caption within the time period, the caption will automatically disappear afterwards. To keep an impact caption, users can simply touch or grab it before it disappears.
Once the caption has interactions with a user, it will no longer be constrained by the living time limitation until users intentionally delete it by holding it and clicking a button on the VR controller (i.e., “X” button for the left-hand controller and “A” button for the right-hand controller).
Additionally, to avoid multiple impact captions overlapping with each other when a user keeps speaking, SpeechCap randomizes the positions for spawning. Specifically, the exact spawn position of a caption is calculated by a fixed position plus random offsets on $x$, $y$, and $z$ dimensions. This strategy creates a “word cloud” effect in front of the speaker’s avatar.
@@ 6.1. Making Social VR Accessible to Deaf and Hard-of-Hearing (DHH) People
Recent research revealed that current VR applications fail to provide sufficient accessibility support for deaf and hard-of-hearing (DHH) people to experience immersive digital content or engage in remote communication and socialization  (Jain et al., 2021; Borna et al., 2024). Captioning systems, widely used to enhance accessibility in other digital media such as videos, remain under-explored in VR contexts  (Kim et al., 2023; de Lacerda Pataca et al., 2023, 2024). SpeechCap not only has similar functionality to previous systems that convert speech into visible captions for DHH users  (Kim et al., 2023; de Lacerda Pataca et al., 2023; Li et al., 2022), but also extends this concept by integrating visual cues and interactivity to convey rich non-verbal information (e.g., valence, excitement).
For example, in a conversation about a vacation plan in social VR with SpeechCap, a DHH user can discern the speaker’s positive mood through bright-colored captions accompanied by a laughing emoji. A shivering motion further indicates the speaker’s excitement about the trip. By dragging and placing impact captions of the names of the must-visit spots mentioned in speech, the speaker can demonstrate the plan with clarity. Additionally, when a speaker picks up a caption labeled “volcano” while discussing hiking and exploration, the corresponding icon reinforces the speech context. Overall, informative captions, engaging visuals, and interactive features in SpeechCap expand opportunities for DHH individuals to connect with others and engage with the virtual world in social VR
@@ 6.2. Enhancing Interactions for Teaching and Learning in Social VR
VR has shown great potential in education by enhancing social presence in remote learning and providing access to learning contexts otherwise unavailable in reality  (Thanyadit et al., 2022; Peng et al., 2022; Jensen and Konradsen, 2018). However, it remains far from mainstream adoption due to challenges in creating easily reviewable educational content, ensuring inclusivity, and supporting collaborative learning  (Jin et al., 2022). In virtual lectures, SpeechCap enables instructors to create instructional cues by placing keyword-based impact captions in the virtual space. Shared interactive captions facilitate interaction between instructors and students, supporting activities like Q&A and hands-on demonstrations.
@@ 6.3. Facilitating Engaging Live Streaming Experiences in Social VR
Recent research suggests that VR streamers face challenges in building emotional connections, as VR headsets obscure facial expressions, making it harder for viewers to perceive their emotions directly  (Wu et al., 2023), while emotional connection is a key factor in live streaming  (Lu et al., 2018). SpeechCap visualizes emotions for streamers by detecting moods in speech and adding appropriate emojis and motion effects to captions, and it also enhances communication by rendering viewers’ reactions as interactive captions in the streamer’s virtual space, fostering engagement. These features of SpeechCap address the need for visible and spatialized objects, creating more engaging live streaming experiences  (Wu et al., 2023; Lu et al., 2018).
@@ 7.1. Participants
We recruited 14 participants (6 female and 8 male, aged between 18 and 35) from local universities through advertisements on social media. None of the participants overlapped with the four experts involved in the previous co-design activity. Five participants (P1 to P5) were novice users who had never used VR head-mounted displays (HMDs) before, while the other nine participants (P6 to P14) self-reported having extensive experience with VR games, social VR, and application development.